home *** CD-ROM | disk | FTP | other *** search
/ NetNews Usenet Archive 1992 #27 / NN_1992_27.iso / spool / sci / space / 16180 < prev    next >
Encoding:
Internet Message Format  |  1992-11-20  |  9.0 KB

  1. Xref: sparky sci.space:16180 sci.astro:12249 alt.sci.planetary:340
  2. Newsgroups: sci.space,sci.astro,alt.sci.planetary
  3. Path: sparky!uunet!elroy.jpl.nasa.gov!kelvin.jpl.nasa.gov!baalke
  4. From: baalke@kelvin.jpl.nasa.gov (Ron Baalke)
  5. Subject: Galileo Update - 11/20/92
  6. Message-ID: <1992Nov20.194126.890@elroy.jpl.nasa.gov>
  7. News-Software: VAX/VMS VNEWS 1.3-4   
  8. Keywords:  Galileo, JPL
  9. Sender: news@elroy.jpl.nasa.gov (Usenet)
  10. Nntp-Posting-Host: kelvin.jpl.nasa.gov
  11. Reply-To: baalke@kelvin.jpl.nasa.gov
  12. Organization: Jet Propulsion Laboratory
  13. Date: Sat, 21 Nov 1992 03:40:35 GMT
  14. Lines: 167
  15.  
  16. Forwarded from Neal Ausman, Galileo Mission Director
  17.  
  18.                                  GALILEO
  19.                       MISSION DIRECTOR STATUS REPORT
  20.                                POST-LAUNCH
  21.                           November 12 - 19, 1992
  22.  
  23. SPACECRAFT
  24.  
  25. 1.  On November 13, the Trajectory Correction Maneuver (TCM-16) was performed
  26. on the spacecraft.  The maneuver used the spacecraft's lateral thrusters to
  27. impart a predicted total delta velocity of 0.89 m/s.  This maneuver was
  28. executed at 1200 bps with the spacecraft pointed approximately 7 degrees off
  29. the Sun.
  30.  
  31.      The spacecraft's performance throughout the activity was nominal. All
  32. RPM (Retro-Propulsion Module) pressures and temperatures and attitude control
  33. indicators were near predicted levels.  After the lateral burn segment, the
  34. sequence planned pointing correction was executed; the sequence planned spin
  35. correction was not needed.
  36.  
  37. 2.  On November 14, after completion of TCM-16, the Energetic Particle Detector
  38. (EPD) instrument was moved from Sector O (TCM safe position) to Sector 5 and
  39. scanning operation initiated to continue science data collection.  The science
  40. scanning mode continued until November 17 at which time the EPD was moved to
  41. Sector 4 and the motor turned off.  The EPD instrument is continuing to
  42. collect science data in Sector 4.
  43.  
  44. 3.  On November 14 and 15, delta Differenced One-Way range (DOR) passes were
  45. performed over DSS-14/63 (Goldstone/Madrid) and DSS-14/43 (Goldstone/Canberra).
  46. Initial results indicate the four delta DOR passes were successfully performed.
  47.  
  48. 4.  On November 16, a periodic RPM 10-Newton thruster maintenance activity was
  49. performed; 10 of the 12 thrusters were "flushed" during the activity.  The
  50. P-thrusters were not flushed because they were used to perform science turn
  51. (SITURN) activities on the same day.  Spacecraft performance throughout the
  52. activity was normal.
  53.  
  54. 5.  On November 16, the spacecraft performed a 13 degree SITURN.  The purpose
  55. of the SITURN was to maintain the spacecraft within plus or minus 10 degrees
  56. of the sun at the current solar distance.  Real-time commands were also sent to
  57. change the downlink telemetry format back to the Low Rate Science (7.68 kbps)
  58. from the Engineering High Rate 1200bps (automatically selected in the maneuver
  59. PA) to collect a unique near-Earth magnetosphere data.
  60.  
  61. 6.  On November 16, the Earth vector and Earth stars were updated to the
  62. spacecraft's current attitude.  This change implements the attitude maintenance
  63. strategy required in the case of an AACS (Attitude and Articulation Control
  64. Subsystem) POR (Power-On Reset).  If a fault occurs which terminates the
  65. currently executing sequence, the spacecraft will maintain its current
  66. attitude.
  67.  
  68. 7.  On November 17, the Relay Radio Hardware (RRH) oscillators No. 1 and 2
  69. were powered on in preparation for the Probe Mission Sequence Test (MST)
  70. scheduled for November 20, 1992.  Shortly after turn on the oscillator
  71. temperatures increased to expected values.
  72.  
  73. 8.  On November 18, real-time commands were sent to turn the Ultraviolet
  74. Spectrometer (UVS) instrument on, load and readout the UVS memory, and start
  75. the UVS microprocessor.  Three memory readouts of the UVS load were completed
  76. and verified prior to turning the UVS microprocessor on. The instrument was
  77. then configured to a nominal state with the master high voltage disabled,
  78. G-channel selected, and the grating stepping.  The UVS is now properly
  79. configured for the start of EE-9 (Earth-Earth 9 sequence). The UVS
  80. supplemental heater was turned off to complete the activity.
  81.  
  82. 9.  On November 19, real-time commands were sent to increase the telemetry
  83. rate from 7.68 kbps to 28.8 kbps in preparation for the Probe Mission Sequence
  84. Test (MST) on the following day.  Additionally, Delayed Action Commands (DACs)
  85. were sent to turn ranging off and two-way non-coherent (TWNC) on prior to the
  86. MST and reconfigure to ranging on and TWNC off after the MST completes.  These
  87. actions were taken to optimize telecommunications performance during the Probe
  88. MST.
  89.  
  90. 10. The AC/DC bus imbalance measurements exhibited some change. The AC
  91. measurement has ranged from 15DN to 17DN and now reads 15DN (3.5 volts).
  92. The DC measurement has ranged from 134DN (15.7 volts) to 142DN (16.7 volts)
  93. and now reads 142DN (16.7 volts).  These measurement variations are consistent
  94. with the model developed by the AC/DC special anomaly team.
  95.  
  96. 11. The Spacecraft status as of November 19, 1992, is as follows:
  97.  
  98.        a)  System Power Margin -  80 watts
  99.        b)  Spin Configuration - Dual-Spin
  100.        c)  Spin Rate/Sensor - 3.15rpm/Star Scanner
  101.        d)  Spacecraft Attitude is approximately 6 degree
  102.            off-sun (leading) and 16 degrees off-earth (lagging)
  103.        e)  Downlink telemetry rate/antenna-28.8kbps (coded)/LGA-1
  104.        f)  General Thermal Control - all temperatures within
  105.            acceptable range
  106.        g)  RPM Tank Pressures - all within acceptable range
  107.        h)  Orbiter Science- PWS, PLS, UVS, EPD, MAG, HIC and DDS
  108.            are powered on
  109.        i)  Probe/RRH - RRH oscillators powered on, temperatures
  110.            within acceptable range
  111.        j)  CMD Loss Timer Setting - 264 hours
  112.            Time To Initiation - 260 hours
  113.  
  114.  
  115. UPLINK GENERATION/COMMAND REVIEW AND APPROVAL:
  116.  
  117. 1.  The EE-12 Final Profile Design Product was approved by the Project on
  118. November 13.  This sequence covers spacecraft activities from January 20, 1993
  119. to January 23, 1993 and from February 8, 1993 to April 12, 1993.
  120.  
  121. 2.  The EE-11 Preliminary Sequence and Command Generation products were
  122. approved by the Project on November 17.  This sequence covers spacecraft
  123. activities from December 7 to December 28, 1992 including the Earth 2 closest
  124. approach on December 8.
  125.  
  126. 3.  The EE-9 Final Sequence and Command Generation package was approved for
  127. transmission by the Project on November 19, 1992.  This sequence covers
  128. spacecraft activities from November 23, 1992 to December 5, 1992.  The
  129. sequence is scheduled to be sent to the spacecraft on November 21.
  130.  
  131. GDS (Ground Data Systems):
  132.  
  133. 1.  A second SPC-60 (Madrid Signal Processing Center) Telemetry Data Flow test
  134. was conducted on Monday, November 16.  The test was successful in establishing
  135. a configuration for the Probe checkouts.  This configuration utilized two
  136. Type A TPA's (Telemetry Processing Assemblies) in parallel, one configured to
  137. the 56K full duplex line and one configured to the 224kbps Simplex line, with
  138. Area Routing Assembly (ARA-1) as Prime and ARA-2 as backup.  The only drop-outs
  139. with this configuration were associated with "line hits" and were recoverable
  140. on ARA ODR (Original Data Recording) replays. The test also successfully
  141. demonstrated the configuration to be used to support 134.4 kbps High Rate
  142. science at Earth 2.  As a result of an excellent effort by DSN (Deep Space
  143. Network) and SPC-60 personnel, data dropouts have been minimal this week.
  144.  
  145.  
  146. TRAJECTORY
  147.  
  148.      As of noon Thursday, November 19, 1992, the Galileo Spacecraft trajectory
  149. status was as follows:
  150.  
  151.     Distance from Earth         14,931,800 km (.10 AU)
  152.     Distance from Sun           161,989,800 km (1.08 AU)
  153.     Heliocentric Speed          118,300 km per hour
  154.     Distance from Jupiter       912,828,500 km
  155.     Round Trip Light Time       1 minute, 38 seconds
  156.  
  157.  
  158. SPECIAL TOPIC
  159.  
  160.  
  161. 1.  As of November 19, 1992, a total of 8684 real-time commands have been
  162. transmitted to Galileo since Launch.  Of these, 3660  were initiated in the
  163. sequence design process and 5024 initiated in the real-time command process.
  164. In the past week, 35 real time commands were transmitted: none were initiated
  165. in the sequence design process and 35 initiated in the real time command
  166. process.  Major command activities this week included commands to reconfigure
  167. the downlink telemetry data rate to 7.68 kbps, turn the UVS instrument on and
  168. configure it for science data collection, turn the UVS supplemental heater off,
  169. switch the telemetry rate to 28.8 kbps, and configure ranging and the TWNC for
  170. the probe MST.
  171.  
  172. 2.  A training rehearsal for the GOPEX experiment operations was successfully
  173. completed November 18th.  The rehearsal was supported by elements of the
  174. Galileo Flight Team (MCT, NAV, and SDT), DSN Ops Team and GOPEX operations
  175. Team at Table Mountain Observatory (TMO) and Kirtland AFB.
  176.      ___    _____     ___
  177.     /_ /|  /____/ \  /_ /|     Ron Baalke         | baalke@kelvin.jpl.nasa.gov
  178.     | | | |  __ \ /| | | |     Jet Propulsion Lab |
  179.  ___| | | | |__) |/  | | |__   M/S 525-3684 Telos | Learn to recognize the
  180. /___| | | |  ___/    | |/__ /| Pasadena, CA 91109 | inconsequential, then 
  181. |_____|/  |_|/       |_____|/                     | ignore it.
  182.  
  183.